# Grassroots Consensus

### v2

@@ Abstract
Grassroots platforms aim to offer an egalitarian alternative to global platforms—centralized/autocratic and decentralized/plutocratic alike. Within the grassroots architecture, consensus is needed to realize platforms that employ digital social contracts, which are like smart contracts except that they are among people not accounts and are executed by these people’s smartphones not by high-performance servers controlled by parties outside to the contract. Key envisioned grassroots platforms include sovereign democratic digital communities and federations, community banks and their grassroots cryptocurrencies, and digital cooperatives.
The grassroots architecture can benefit from a consensus protocol that is (i) quiescent, (ii) efficient during both low- and high-throughput, (iii) responsive, (iv) blocklace-based, (v) UDP-ready, and (vi) grassroots. The Grassroots Consensus protocol addresses all these requirements while having competitive performance in both low- and high-throughput scenarios and being one of the most concise and elegant consensus protocols for partial synchrony. It achieves that by building on two cutting-edge consensus protocols—the quiescent high-performance Morpheus and the blocklace-based Cordial Miners, improving the latter’s dissemination protocol and making it UDP-ready, and extending the protocol with a constitution and a constitutional amendment component, making it grassroots.
@@ 1 Introduction
Grassroots. Grassroots platforms [44, 42, 40, 45, 46] aim to offer an egalitarian alternative to global platforms—centralized/autocratic and decentralized/plutocratic alike. A grassroots platform is different from a global platform in that it may have multiple instances that can operate independently of each other, without coordination or reliance on any third-party or global resource other than the network itself; furthermore, different instances of the same grassroots platform may interoperate and coalesce into ever-larger instances, possibly, but not necessarily, resulting in a single global instance.
Decentralized platforms for global cryptocurrencies [31, 52] and smart contracts [15, 9, 30] supporting Decentralized Finance (DeFi) [36, 21] and Decentralized Autonomous Organizations (DAOs) [19, 16, 49] often employ a consensus protocol as their foundation. Hence the surge in interest in high-performance and high-throughput consensus protocols [53, 23, 24, 27, 5, 22].
Grassroots platforms, on the other hand, are not consensus-based. Their foundational platform consists of a grassroots social graph implemented via a blocklace—a DAG-like generalization of the linear blockchain that is a universal Conflict-free Replicated Data Type (CRDT) [3]. The grassroots architecture [42, 44] facilitates the implementation of a grassroots social network [41] and grassroots cryptocurrencies [43, 27], which are also consensus-free, on top of the grassroots social graph. So where does consensus come into play?
Consensus. Within the grassroots architecture, consensus is needed to realize higher-level platforms that employ digital social contracts—the grassroots counterpart of smart contracts. Briefly, while a smart contract [15] is among anonymous accounts and runs on a consensus protocol executed by third-party servers, a social contract is among people known to each other and runs on a consensus protocol executed by the smartphones of these very same people [44, 46]. The two types of consensus-based contracts are compared in Table 1.
Table 1: Comparing Consensus-Based Smart Contracts and Social Contracts
@@
Example: Grassroots Federation. A key application of digital social contracts mentioned in Table 1 is the grassroots federation of digital communities. Grassroots Federation [46] aims to address the egalitarian formation and the fair democratic governance of large-scale, decentralized, sovereign digital communities, the size of the EU, the US, existing social networks, and even humanity at large. A grassroots federation evolves via the grassroots formation of digital communities and their consensual federation. Such digital communities may form according to geography, jurisdiction, affiliations, relations, interests, causes, and more. Small communities (say up to $100$ members) govern themselves; larger communities—no matter how large—are governed by a similarly-small assembly elected by sortition [20, 46] among its members. It is expected that the constitutional democratic governance of each community would be specified by a digital social contract running on top of a grassroots consensus protocol, which in turn would be executed by the members of said small community or its assembly, one protocol instance per community. Such a consensus protocol should satisfy all the requirements explained next.
Requirements of Grassroots Consensus. Small communities engaged in democratic conduct, small community banks, small cooperatives, etc., would have a transaction rate much lower than network latency. Hence they would need a (i) quiescent protocol in which participants do nothing (in particular, send no messages) until there is a new transaction, and when subsequent transactions are issued at low-throughput, finalize each quickly and efficiently. As during low-throughput transactions are issued one-at-a-time, there is no symmetry to break or conflicts to resolve and hence—ideally—the agent issuing a transaction should be able to finalize it without the help of a leader, namely during low-throughput the protocol should be leaderless.
At the same time, the protocol should also work efficiently and with low-latency on smartphones in a high-throughput scenario. Namely, it should be an (ii) efficient consensus protocol during both low- and high-throughput. In addition, a standard requirement, termed (iii) responsiveness, is that in the good case the protocol operates at network speed, with performance unaffected by an overly-conservative estimation $\Delta$ of the actual least upper bound on network delay $\delta$.
A fourth requirement relates to the blocklace: As the lower-level grassroots platforms (Social Graph and Social Networks [41], Cryptocurrencies [43, 27]) are blocklace-based, it would be very beneficial and integrable to also have a (iv) blocklace-based consensus protocol. This is not a limitation as there are plenty of efficient and high-throughout blocklace-based protocols already, starting from Cordial Miners [24] and its extensions and improvements [5, 22]. We note that a blocklace [3] is not “just a DAG” the same way a blockchain is not “just a sequence”. In a blocklace, directed edges are realized by signed cryptographic hash pointers: A $p$-block (block by agent $p$) $b$ includes a cryptographic hash of the rest of $b$, signed by $p$, which serves as its unique (whp) identifier, also referred to as a pointer to $b$. An edge from the $p$-block $b$ to the $q$-block $b^{\prime}$ is realized by $b$ including a pointer to $b^{\prime}$. Thus, similarly to a blockchain, the blocklace is:
@@
    Tamper-proof: A bit in a block cannot be changed without this being detected;
@@
    Non-repudiable: The creator of a block cannot deny having created it;
@@
    Acyclic: Even Byzantine agents cannot create blocks that form/close a cycle.
However, unlike the blockchain, which is extended competitively, with the point-of-contention being adding the next block to the most-recent block, a blocklace is a naturally-cooperative data type, as agents may add blocks to the blocklace at will without conflicting with each other. In fact, the blocklace is a Universal Conflict-free Replicated Data Type (CRDT) [3].
The grassroots architecture in general, and the consensus protocol presented here in particular, are geared for smartphone-based execution. Namely, the intention is that the parties to a social contract will run both the contract and its underlying consensus protocol on their smartphones. The vagaries of the Internet—NATs and firewalls—and the fact that mobile phones change their IP address as they roam about, make phone-to-phone communication difficult. In particular, opening a direct TCP connection between two phones is either impossible, or impossible without the help of a proxy server, and even if established, it has to be re-established whenever one of the phones changes their IP address. Hence, UDP is a much preferable protocol for phone-to-phone communication, and is the one typically used for audio and video phone-to-phone conversations. Thus, the fifth requirement is that the protocol work not only on reliable networks, as typically assumed by consensus protocols, but also on unreliable networks namely be (v) UDP-ready.
The final sixth requirement for a protocol for grassroots platforms is, naturally, that the protocol be (vi) grassroots. Informally, this means two things: (i) That a group of agents $P$ running the protocol can do so unscathed if embedded within a larger group of agents $P^{\prime}\supset P$. Any permissioned consensus protocol satisfies that. (ii) That the protocol can do more if $P$ is embedded within a larger group of agents $P^{\prime}$.
Any platform that operates on a shared global resource or employs a global replicated (Blockchain [31]), or distributed (IPFS [7], DHT [35]) shared data structure, or distributed pub/sub systems with a global directory [13, 14, 8], are all not grassroots. Server-based federated systems such as Mastodon [34] are also not grassroots, informally due to the control exerted by each server-operator on its members; formally, since a group of people of any size cannot function without first connecting to such a server.
For a permissioned consensus protocol to be grassroots, the set of agents $P$ needs to be “reconfigured” to include agents in $P^{\prime}\setminus P$. The question of reconfiguration has been studied extensively for decades [26, 17, 2, 47]. While the original informal exposition of the problem in the Paxos paper was with respect to a self-governed parliament [26], subsequent works in the context of permissioned consensus assume that the initial ‘permissioned’ agents are determined by an external authority, and so is their ‘reconfiguration’. Proof-of-Stake protocols such as Ethereum 2.0 offer a stake-based approach to reconfigure the participating agents on every epoch [25, 10].
A grassroots system cannot have external authorities and hence protocol reconfiguration must be entirely under the control of the very same agents that execute the protocol, as envisioned in the Paxos paper [26]. Furthermore, to be egalitarian rather than plutocratic, agents should have equal power in determining reconfiguration decisions. Specifically, in the grassroots context the challenge relates to the consensus protocol being executed by a self-governed community (the parties to the social contract running on top of the consensus protocol), allowing them to add or remove members at will, as well decide to change other aspects of its operation, for example the supermajority required to finalize a decision. To address that, we extend the protocol with a constitution that specifies the agents that execute the protocol as well as two key protocol parameters—the supermajority $\sigma$ by which decisions, including constitutional amendment decisions, are binding, and a time $\Delta$ estimating the unknown least upper bound on network delay during synchrony, $\delta$.
To summarize, the grassroots architecture can benefit from a consensus protocol that is (i) quiescent, (ii) efficient during low- and high-throughput, (iii) responsive, (iv) blocklace-based, (v) UDP-ready, and (vi) grassroots. The Grassroots Consensus protocol was conceived to address all these requirements.
Paper structure. Section 2 presents the Grassroots Consensus protocol assuming a reliable network, as well as a prevailing constitution $(P,\sigma,\Delta)$, that specifies the set of agents $P$, the supermajority $\sigma$ among the agents by which the protocol operates, and a presumed upper-bound $\Delta$ on network delay during synchrony. Section 3 proves the safety and liveness of Grassroots Consensus. Section 4 extends the protocol for eventually-reliable networks, making it UDP-ready. Section 5 proves that the Grassroots Consensus protocol is indeed grassroots. Section 6 extends Grassroots Consensus with constitutional amendment, with which the agents may amend these three protocol parameters, and present democratic processes geared for amending each of the three parameters. Section 7 concludes.
2 Grassroots Consensus

### v4

# Constitutional Consensus

@@ Abstract
Consider people with smartphones operating without external authorities or global resources other than the network itself. In this setting, high-end applications supporting sovereign democratic digital communities, community banks, and digital cooperatives require consensus executed by community members, which must be reconfigurable to support community dynamics.
The Constitutional Consensus protocol aims to address this need by introducing constitutional self-governance to consensus: participants dynamically amend the participant set, supermajority threshold, and timeout parameter through the consensus protocol itself. We achieve this by enhancing a DAG-based protocol (like Cordial Miners) with participant-controlled reconfiguration, while also supporting both high- and low-throughput operation (like Morpheus), remaining quiescent when idle. This three-way synthesis uniquely combines: (1) constitutional amendments for self-governance, (2) a cryptographic DAG structure for simplicity, parallelism, and throughput, and (3) both high- and low-throughput operation. The protocol achieves consensus in $3\delta$, maintains $O(n)$ amortized communication complexity during high throughput, and seamlessly transitions between modes. The basic protocol (without constitutional amendments) realizes these features in 25 lines of pseudocode, making it one of the most concise consensus protocols for eventual synchrony.
@@ 1 Introduction
Participant-Controlled Consensus. Consider people who wish to form self-governing digital communities without relying on servers or external authorities. These participants—whether members of community banks, digital cooperatives, or democratic assemblies—need consensus protocols they can execute on their smartphones while maintaining complete control over membership and governance rules. Unlike traditional reconfiguration where external authorities determine changes [40, 2], these communities require constitutional self-governance: participants must collectively control who joins, how decisions are made, and how the protocol operates, all through the consensus protocol itself. This paper presents Constitutional Consensus, a protocol that enables such self-governance by allowing participants to dynamically amend the participant set $P$, supermajority threshold $\sigma$, and timeout parameter $\Delta$ through democratic decisions implemented within the protocol.
Digital Social Contracts. These self-governing communities implement digital social contracts [9]—deterministic state machines that encode agreements among people who know each other. Like smart contracts, they provide computational guarantees about contract execution. However, while smart contracts [10] operate among anonymous accounts and run on consensus protocols executed by third-party servers, social contracts are among identified people and run on consensus protocols executed by the participants’ own devices. Existing consensus-based smart contracts, notably in Decentralized Finance (DeFi) platforms [10, 7, 25], have driven development of high-performance consensus protocols [43, 17, 18, 4, 16], but these operate among pseudonymous accounts on global networks rather than identified people in local communities. This fundamental difference enables democratic rather than plutocratic governance, as shown in Table 1.
Table 1: Comparing Smart Contracts and Consensus-Based Digital Social Contracts
@@
Governance: Token-weighted voting (plutocracy) One person, one vote (democracy)
Amendment Authority: External governance tokens or DAOs Participants themselves via supermajority
Parameter Control: Fixed in code or controlled by token holders Democratically adjustable by participants
We note that while smart contracts are necessarily consensus-based, digital social contracts are not. Some grassroots platforms, notably grassroots social networks [35] and grassroots cryptocurrencies [37, 22], do not require consensus on the total ordering of the acts of their participants: Their safety depends only on the agents obeying the social contract [38]. This can be ensured by mechanisms weaker then consensus, e.g. mutual attestations among agents operating within Trusted Execution Environments (TEEs) [30]. However, applications requiring collective agreement—such as community banks with multiple signatories or democratic governance with voting—do require consensus to ensure all participants agree on the order of transactions and decisions.
Technical Requirements. Communities implementing digital social contracts face unique technical challenges. Small communities may have transaction rates much lower than network latency, requiring a protocol that remains quiescent—sending no messages when idle. Yet the same protocol must efficiently handle high-throughput periods when the community is larger and/or more active. During low throughput, transactions arrive one-at-a-time with no conflicts to resolve, so ideally any participant should be able to finalize their transaction without designated leaders. During high throughput, the protocol should match the performance of modern consensus protocols.
Specifically, an effective protocol for self-governing communities requires:
@@
    Quiescent operation: No activity when idle, spontaneous leader emergence for isolated transactions
@@
    Dual-mode efficiency: Seamless transition between low and high throughput modes
@@
    Responsiveness: Network-speed operation unaffected by conservative timeout estimates
  4. 4.
    DAG-based structure: Leveraging blocklace [3] parallelism and integration
  5. 5.
    Constitutional amendments: Participant-controlled governance of protocol parameters
Technical Foundation. Constitutional Consensus achieves these requirements through a three-way synthesis. From Morpheus [23], it inherits dual-mode operation with spontaneous leaders during low throughput and formal leadership during high throughput. From Cordial Miners [18], it adopts the blocklace structure for parallelism and efficiency. To these, it adds constitutional amendment mechanisms enabling true self-governance. The protocol operates in waves of three rounds. During low throughput, any participant with a transaction can spontaneously lead a wave, achieving finality in $3\delta$. During high throughput, formal leaders are selected round-robin, maintaining $O(n)$ amortized communication complexity. The basic protocol (without amendments) requires only 25 lines of pseudocode, making it one of the most concise consensus protocols for eventual synchrony.
Constitutional Self-Governance. The key innovation lies in how participants control their own governance. Unlike smart contracts that rely on external governance tokens or DAOs [12], constitutional amendments are decided by the participants themselves through supermajority vote. While Lamport’s original Paxos envisioned a self-governing parliament [21], subsequent reconfiguration protocols [42, 2, 40, 20] assume external authorities determine membership changes. Constitutional Consensus returns to Paxos’s original vision of true self-governance. Unlike Proof-of-Stake systems where validators must stake capital [8], participation requires only the democratic approval of existing members. This creates a fundamental architectural difference: while blockchain systems typically support only a single global instance whose security depends on token value, Constitutional Consensus enables unlimited independent communities to self-govern simultaneously.
Connection to Grassroots Architecture. This capability for multiple independent instances operating without central coordination or global resources makes Constitutional Consensus inherently grassroots [38]. A community in Bangladesh, a cooperative in Brazil, and a federation in Berlin can each run their own instance, potentially with overlapping members, without any coordination or competition for resources. This architecture—multiple local instances rather than a single global instance—enables the broader grassroots vision of egalitarian digital platforms that operate without central authorities [36, 34]. Within this grassroots architecture, Constitutional Consensus provides the consensus layer for applications requiring collective agreement, complementing consensus-free grassroots platforms like social networks [35] and cryptocurrencies [37, 22]. Together, these protocols enable communities to build sovereign democratic digital infrastructure from the ground up, controlled entirely by their members rather than external authorities or token holders. This distinguishes Constitutional Consensus from global platforms—whether centralized like Facebook or decentralized like Bitcoin [26] and IPFS [5]—that can support only a single instance.
Example: Grassroots Communities, Banks, and Federations. As a concrete scenario for grassroots platforms, consider a small town whose residents wish to form a digital community. They begin by deploying independent instances of consensus-free grassroots platforms on their smartphones—creating their own grassroots social network [35] and trading via self-minted grassroots cryptocurrencies [37, 22]. Crucially, they do this without coordination with any global network or central authority. When their needs evolve to require collective agreement—community banks with multiple signatories, constitutional governance, or federation with neighbouring towns—they deploy their own Constitutional Consensus instance. This instance operates independently yet can interoperate with other communities’ instances, embodying the grassroots principle of multiple local instances rather than a single global system. Through Grassroots Federation [39], these autonomous instances can scale democratically from local to potentially global governance. Small communities ($\approx 100$ members) govern themselves directly; larger ones use sortition-selected assemblies [14, 39]. Each community’s governance runs as a digital social contract on their own Constitutional Consensus instance, executed by members on their smartphones—no external servers or authorities required.
Model. As deterministic fault-tolerant consensus is not solvable in asynchronous systems [13], we assume an eventual synchrony model, where communication delays are at most $\delta$ after some Global Stabilization Time (GST) that is unknown to the agents. We consider Byzantine failures, as some malicious actors may infiltrate the network, but we assume a bounded ratio of Byzantine agents (to be specified later), since, by and large, the agents trust each other.
Paper structure. Section 2 provides the model and problem definitions. Section 3 presents a single-epoch Constitutional Consensus protocol, with only an initial constitution and no constitutional amendments. Section 4 proves the safety and liveness of single-epoch Constitutional Consensus. Section 5 adds support for constitutional amendments – for each of the three parameters $P,\sigma$, and $\Delta$, we present a democratic process for amending it. We then present the protocol extension needed to realize constitutional amendments and prove its safety and liveness. Section 6 proves that Constitutional Consensus is grassroots. Section 7 concludes.
2 Model and Problem
Model. We assume a set of agents $\Pi$, each endowed with a unique key-pair and identified by its public key $p\in\Pi$. While $\Pi$ is potentially-infinite, we refer only to finite subsets $P\subset\Pi$ of it. A correct agent follows the protocol. An agent that does not follow the protocol is faulty or Byzantine.
We assume an eventual synchrony (aka partial synchrony) model in which every message sent between correct agents is eventually received. Message latency is unbounded, but in every run there is a time called the Global Stabilization Time (GST), which is unknown to the agents, after which all messages arrive within a least upper bound $\delta$ time for some $\delta>0$.
Note that the abstraction of reliable messaging does not mean that agents must communicate via point-to-point TCP connections. Rather, it abstracts away a dissemination layer that can be implemented using acks/nacks, gossip, network coding, etc., and is orthogonal to the consensus protocol in this paper.
A protocol proceeds through a sequence of epochs, each commences with a new constitution (defined next), referred to as the prevailing constitution during the epoch, and ends when the constitution is amended, namely replaced with a new constitution. The constitution defines the set of participants in the epoch, as well as key protocol parameters:
Definition 2.1 (Constitution, $\sigma$-Supermajority).
A constitution is a triple $(P,\sigma,\Delta)$ where:
  1. 1.
    $P\subset\Pi$ is a set of agents, with $n:=|P|$,
  2. 2.
    $\frac{1}{2}\leq\sigma<1$ is a fraction that specifies a supermajority among the agents $P$,
  3. 3.
    $\Delta>0$ is a presumed upper-bound on message delay after GST among $P$.
We may omit $\Delta$ when its value is immaterial (e.g., when discussing protocol safety) and refer to the constitution $(P,\sigma)$. A $\sigma$-supermajority among $P$ is a fraction $Q\subseteq P$ such that $|Q|>\sigma n$.
Supermajorities and faulty agents. Typically, a permissioned consensus protocol assumes that there are at most $f<n$ faulty agents among $P$, and with this assumption we define $\sigma:=\frac{n+f}{2n}$. Equivalently, $\sigma$ is chosen with the assumption that there are at most $f=(2\sigma-1)n$ faulty agents in $P$. For example, for $f=0$, $\sigma=\frac{1}{2}$; for $f=\frac{1}{2}n$, $\sigma=\frac{3}{4}$; and if everyone but one agent is faulty then $\sigma=\frac{2n-1}{2n}=1-\frac{1}{2n}$, namely unanimity is required. For $n=3f+1$ (the standard assumption), $\sigma=\frac{n+\frac{n-1}{3}}{2n}=\frac{2}{3}-\frac{1}{6n}$, namely a supermajority greater-or-equal to $\frac{2}{3}n$ is required.
Typically, permissioned consensus protocols wait to hear from “at least nfn-f agents among PP”, which in the standard case of $n=3f+1$ (but not otherwise) happens to be equivalent to “a $\sigma$-supermajority among PP”.
Constitution safety and liveness. A constitution thus defines the protocol’s assumptions about the environment: $P$ is the set of agents running the protocol; there is a correct $\sigma$-supermajority in $P$ and every $\sigma$-supermajority in $P$ includes a majority of correct agents, hence the intersection of any two $\sigma$-supermajorities contains a correct agent; $\Delta\geq\delta$ is an upper bound on the communication time after GST.
Definition 2.2 (Safe and Live Constitution).
A constitution $(P,\sigma,\Delta)$ is safe if $P$ includes at most $f=(2\sigma-1)n$ faulty agents, and live if after GST any message sent among correct agents in $P$ arrives within $\Delta$.
Constitutional Consensus problem definition. A constitutional consensus protocol may have multiple instances. Each instance commences with an initial constitution $(P,\sigma,\Delta)$ that defines its participating agents $P$ and the key protocol parameters $\sigma$ and $\Delta$. As in standard consensus, participants spontaneously issue transactions (these are the inputs to the protocol) and every agent outputs a sequence of transactions.
In addition, constitutional consensus assumes an that participants engage in an unspecified democratic decision process that may result in a decision $d$ to amend the prevailing constitution. The discussion of such democratic processes is deferred to Section 5. For the purpose of the constitutional consensus protocol, all is needed is to assume: (i) the existence of such a process that may result in constitutional amendment decisions, which become known to the participants once made; (ii) that participants can verify a constitutional amendment decisions $d$ to be valid according to the rules specified in Section 5. A valid constitutional amendment decision $d$ defines a new constitution $(P^{\prime},\sigma^{\prime},\Delta^{\prime})$, in which the participants $P^{\prime}$ and/or the parameters $\sigma^{\prime}$ and $\Delta^{\prime}$ are the result of amending the prevailing constitution.
In the following, constitution refers to the new constitution of a valid constitutional amendment decision, and an initial constitution is the new constitution of a valid amendment of a constitution with an empty population. Valid amendments are provided to the consensus protocol as inputs by participants that know of them. Thus, the output sequence of a constitutional consensus protocol includes both transactions and valid constitutional amendment decisions, as follows:
Definition 2.3 (Constitutional sequence, epoch, epoch sequence, prevailing constitution, participants).
A constitutional sequence is a sequence $T=t_{1},t_{2},\dots$ of transactions and valid constitutional amendment decisions beginning with an initial constitution. Each constitutional amendment decision $d$ in the sequence defines a new epoch, its epoch sequence is a maximal contiguous subsequence of $T$ that begins with $d$ and does not include another such decision, and its prevailing constitution is the new constitution of $d$. An agent participates in an epoch if it is a member of the population of its prevailing constitution.
The constitutional output sequences produced by correct agents executing a constitutional consensus protocol instance must satisfy the following safety and liveness conditions:
Definition 2.4 (Constitutional Consensus Correctness).
In any run of an instance of constitutional consensus in which all valid constitutional amendment decisions made are safe and live, there exists a constitutional sequence $T=t_{1},t_{2},\dots$ for which the following safety and liveness properties hold:
  1. 1.
    Safety:
    Consistency
    Every correct agent’s output sequence is a subsequence of $T$.
    Validity
    If a correct agent $p$ outputs a transaction $t$, then $p$ is a participant in $t$’s epoch in $T$.
  2. 2.
    Liveness:
    Epoch Liveness
    If a correct agent $p$ is a participant in an epoch $d$ in $T$, then $p$’s output sequence eventually includes all the transactions in $d$’s epoch sequence. Furthermore, if $d$ is the last epoch in $T$, then every transaction input by a correct participant in the epoch $d$ is included in $T$.
    Constitution Amendment Liveness
    If a valid constitutional amendment decision $d$ is provided as input to the protocol by a correct agent, then $T$ eventually includes $d$.
Note that if there is no decision to amend the initial constitution then the protocol run has a single epoch with a fixed set of participants. In this case, the definition above coincides with the standard consensus correctness definition.
Note also that progress is guaranteed (after GST) even if the constitution changes infinitely many times. However, normal transactions might starve due to frequent amendments. This definition allows the implementation to prioritize constitution changes over regular transactions, allowing the latter to starve. In practice, amendments are expected to be rare compared to transactions, so we believe this modus operandi to be reasonable.
3 Single Epoch Constitutional Consensus
Here we present the behaviour of the Constitutional Consensus protocol during a single epoch. The protocol is based on the blocklace data structure, which we define in Section 3.1. The protocol’s concepts are detailed in Section 3.2 and its pseudo-code in Section 3.3.